The sum of three consecutive integers is 108. What are the integers?
Answer by ewatrrr(24785)

Hi,
Let n be the smallest of the three consectuive intergers.
Then (n+1) and (n+2) would be the other two
.
question states*
n +(n+1) + (n+2) = 108
.
3n + 3 = 108
.
3n = 105
n = 35
.
35, 36, 37 are the integers
.
check your answer
35 + 36 + 37 =108